A speed boat heads west at 101km/h for 29.0 min. it then travels at 60.0 south of west at 90.0 km/h for 12.0 What is the average speed and velocity of trip?
Answer #1

total didtance travelled d = 101*29/60 + 90*12/60

= 66.817 km

avarage seed = total didtancetravlleed/time taken

= 66.817*60/(29+12)

= 97.8 km/h
Answer #2

total displacemnt, d = sqrt( (101*29/60)^2 + (90*12/60)^2 + 2*(101*29/60)*(90*12/60)*cos(45))

= 62.85 km

Average speed = dispalacement/time taken

= 62.85*60/(29+12)

= 91.97 km/h